4 (2.52). The minimum stopping distance for a car traveling at a speed of 30 m/s is 60 m, including the distance traveled during the drivers reaction time of 0.50 s. a) What is the deceleration of the car while braking? b) What is the minimum stopping distance for the same car traveling at a speed of 40 m/s with the same reaction time and deceleration?
Answer #1

a)

Using the equation of motion :

v^2 = u^2 + 2as

So,

0^2 = 30^2 + 2*(-a)*(60 - 30*0.5)

So, a = 10 m/s2 <---------- answer (a)

b)

v^2 = u^2 + 2as

So, 0^2 = 40^2 + 2*(-10)*s

So, s = 80 m

So, minimum stopping distance , d = 80 + 40*0.5 = 100 m <------ answer